### v1

# Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector

Abstract
The Circular Electron-Positron Collider (CEPC), a proposed next-generation $e^{+}e^{-}$ collider to enable high-precision studies of the Higgs boson and potential new physics, imposes rigorous demands on detector technologies, particularly the vertex detector. JadePix-3, a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector, targets a spatial resolution of 3 $\rm{\upmu m}$ to meet the key requirement. This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and detection efficiency. Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate. Simultaneously, an extensive methodology for MAPS characterization is established, which determined the optimal operational working conditions for the JadePix-3 sensor. Furthermore, these findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
keywords:
CEPC , Monolithic active pixel sensors , CMOS , Particle detection
journal: NIM-A
[…]
1 Introduction
The discovery of the Higgs boson has spurred interest in circular $e^{+}e^{-}$ colliders as premier facilities for high-precision studies of the Standard Model Higgs boson and other fundamental particles. Planned designs aim to cover a wide range of center-of-mass energies—from the Z pole (91 GeV) and WW threshold (161 GeV) to the Higgs factory region (240 GeV) and the $t\bar{t}$ threshold (340–365 GeV)—using an $e^{+}e^{-}$ accelerator with a circumference of approximately 100 km. In addition, the same tunnel could later accommodate a hadron collider with a center-of-mass energy nearly 100 TeV. The Circular Electron-Positron Collider (CEPC) [1, 2] has been proposed to pursue this road map, envisioned to investigate the Higgs boson further and reveal new physics potentially.
CEPC’s ambitious physics objectives impose stringent performance requirements on its detectors—particularly the vertex detector. As outlined in Table 1, the CEPC vertex detector must achieve exceptional spatial resolution, maintain a low material budget, minimize power consumption, handle high hit densities, and exhibit robust radiation tolerance to cope with the beam-induced background near the interaction point. In this context, only Monolithic Active Pixel Sensors (MAPS) offer a promising solution to meet these demanding criteria [3, 4].
Therefore, considerable R&D efforts have been undertaken to develop advanced pixel sensor prototypes with monolithic technology for the CEPC vertex detector, such as the TaichuPix series [5, 6, 7] and the JadePix series. The JadePix series is specifically developed to meet the CEPC vertex detector’s demanding criteria, offering high position resolution along with enhanced efficiency. Building on the progressive improvements of its predecessors—JadePix-1, which was to verify the sensor fabrication and optimize the pixel architecture [8, 9], and JadePix-2, which featured an improved analog front‐end design [10]—JadePix-3 represents the latest development and aims for a position resolution of 3 $\rm{\upmu m}$ [11]. The next part presents the detailed design specifications of JadePix-3.
Table 1: Demands for the CEPC vertex detector [1].
Single point resolution $\rm{\upmu m}$
Thickness of sensor chips 50 $\rm{\upmu m}$
Power consumption <50 $\rm mW/cm^{2}$
Timestamp precision $\rm{\upmu s}$
Pixel-hit rate 40 $\rm MHz/cm^{2}$
Radiation tolerance TID, 3.4 Mrad/year
NIEL, $\rm 6.2\times 10^{12}n_{eq}/(cm^{2}year)$
Figure 1: (a) Layout of the entire sensor mask; (b) Circuit topology for the analog front-end [11].
1.1 Overview of JadePix-3 design
The JadePix-3 is designed for a high spatial resolution of 3 $\rm{\upmu m}$. With a pixel matrix consisting of 512 rows and 192 columns, the sensor covers a sensitive area of 8.19 $\times$ 4.85 $\rm{mm^{2}}$ as shown in Fig. 1. The sensor is divided into four distinguishable sectors in the column direction. As presented in Tab. 2, while the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3. The pixel arrangement in each sector is mirrored, with shared vertical metal lines for biasing and control signals.
Fig. 1 demonstrates JadePix-3’s front-end circuit design. As described in Ref. [11] in detailed, the initial amplification stage utilizes a cascode amplifier, which provides high gain while minimizing noise, with its output referred to as OUT_A. Subsequently, the amplified signal is fed into a discriminator stage, where it is compared against a predefined threshold. The resulting digital output, designated as OUT_D, enables efficient hit detection.
Figure 2: Schematic diagram of the cross-section of the sensor when a reverse bias is applied to the P-substrate via a P-well.
Table 2: Analog front-ends and digital logic blocks for different sectors.
Sector pixel layout digital logic block Remarks
0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 SR flipflop
1 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V1 D flipflop, w/o MASK
2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 SR flipflop, w/o MASK
3 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 Increased current for the front-end
The chip employs both rolling shutter (RS) and global shutter (GS) readout sequence modes. In the conventional rolling shutter mode, one row is activated for reading at a time [12]. This process requires 192 $\rm ns$ to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$. While in global shutter, all 512 rows are frozen for a period of time during readout.
The JadePix-3’s sensor design is shown in Fig. 2. A reverse bias $V_{\rm sub}$ can be applied on P-well and P-substrate, expanding the depletion region around the N-well electrode.
1.2 Characterization Motivation
Previous initial characterization proved that JadePix-3 had achieved the general design requirements under zero substrate reverse bias. The result revealed minimum thresholds ranging from 90 to 140 electrons, with a fake-hit rate below $\rm 1\times 10^{-10}$ per frame per pixel [11]. In the beam test for the JadePix-3 telescope, the sensor showed spatial resolutions of 5.2 $\rm{\upmu m}$ and 4.6 $\rm{\upmu m}$ in two dimensions, with a maximum detection efficiency approaching 99.0%  [13]. However, the full performance potential of the JadePix-3 sensor remains incompletely characterized, particularly concerning the effects of reverse bias.
The performance of MAPS is significantly influenced by the bias voltage applied to the substrate. This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [14, 15]. On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry tends to increase due to the body effect, consequently altering the circuit’s operating status. Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential. These metrics are vital for the R&D of the CEPC vertex detector.
Consequently, this work provides a thorough study of reverse bias effects, which is imperative for optimizing sensor design and ensuring its suitability for high-precision particle tracking applications.
This work also introduces an integrated and systematic test methodology (detailed in Section 2), which was established to guide the subsequent comprehensive, lab-based characterization of MAPS behavior under varying reverse bias conditions. While most studies hold the substrate bias fixed or explore a narrow range [16, 17], our approach systematically sweeps the reverse bias from 0 V to –6 V under identical threshold settings. This methodology establishes comparable operating conditions across different bias voltages, thereby facilitating an unambiguous assessment of performance changes predominantly attributable to the bias variation itself.
2 Workflow for Comprehensive Characterization
This work employs a structured, integrated methodology (see Fig. 3):
  • 1.
    Phase 1: Measurement Setup, which is introduced in Section 3, focuses on the preparation stage and establishes a stable experimental setup that minimizes mechanical and environmental perturbations, thereby ensuring reliable test conditions.
  • 2.
    Phase 2: Working-Point Tuning is introduced in Section 4. This phase establishes a series of optimal operating points through a two-stage procedure.
    1. (a)
      Target threshold: Target working-points are defined as distinct operating configurations, one for each applied bias level, that all share a consistent average threshold. This is achieved by applying reverse bias and meticulously tuning DAC parameters to meet the predefined target threshold.
    2. (b)
      ITHR scan: Subsequently, an ITHR scan is performed from each of these target working-points to generate an expanded set of conditions, offering a range of distinct threshold levels for each bias.
    This tuning not only confirms the feasibility of operating under varying reverse biases but also lays the groundwork for all subsequent measurements.
  • 3.
    In Phase 3: Comprehensive Performance Evaluation, which is illustrated in Section 5, three sequentially interlinked investigations were carefully designed and performed:
    1. (a)
      Sensor Electrical Properties Characterization: Detailed measurements of leakage current and input capacitance—both regarded as critical sensor performance metrics—were conducted to understand the intrinsic electrical behavior of the sensor under bias.
    2. (b)
      Noise Performance Analysis: Fake-hit rate is measured to assess noise characteristics, since maintaining a low fake-hit rate is critical for ensuring the effectiveness of hit data for the subsequent measurements.
    3. (c)
      Detection Response Evaluation: The sensor’s response is further assessed by exposing it to external sources, including radioactive sources and an infrared laser. The laser tests, in particular, provide high-resolution mapping of in-pixel charge collection efficiency and serve as a proxy for evaluating sensor performance in response to external stimuli.
Figure 3: Detailed workflow to comprehensively characterize the MAPS sensor.
This integrated workflow not only delivers a robust, reproducible characterization of JadePix-3 in the laboratory—fully capturing how reverse bias shapes every aspect of its response—but also provides a transferable framework for future MAPS R&D.
3 Experimental Setup
The JadePix-3 chip has been wire-bonded to an adaptor board and then connected to an external FPGA-based DAQ system (see Fig. 4). The FPGA board configures the chip parameters and transfers data to the PC via Ethernet. A shock-absorbing platform was employed to maintain system stability and prevent displacement of the test source, either a radioactive source or a laser source, and the sensor during the measurements. A source meter is used to apply reverse bias ranging from 0 to -6 V to the substrate.
Figure 4: (a) Schematic view of the test system; (b) Setup of the infrared laser test system. All devices are fixed or placed on an optical table.
To proceed laser experiment under the demand of high spatial resolution, a laser test system featuring three high-precision motor controller is deployed (see Fig. 4). The spatial position in three orthogonal directions of the laser source is regulated precisely by these three motors, each has a repeatable position resolution of 1 $\rm{\upmu m}$ and a minimum step size of 0.1 $\rm{\upmu m}$. Details about the experiment parameters and test method will be discussed thoroughly in Section 5.5.
4 Working-point Selection
This section details the methodology for selecting operational working-points of the JadePix-3 sensor. The primary objective is to establish uniform thresholds that enable the comparison of the sensor’s performance across varying substrate reverse biases. Central to this process is the accurate measurement of pixel thresholds, achieved using the S-curve method.
This methodology involves giving each pixel a certain number of electrical pulses and counting the hits while increasing the injected charge, which is regulated as $V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor that has a designed value of $0.18~{}{\rm fF}$ marked in Fig. 1, and $V_{\rm pulse}$ is the external voltage for test-pulse injection.
As the injected charge approaches the pixel threshold, the hit count rises along an S-shaped curve, typically modeled by fitting an error function. The pixel threshold is then defined as the injected charge level at which this fitted error function reaches half of its maximum value. Threshold variations between pixels are quantified by the fixed pattern noise (FPN), defined as the standard deviation of the mean threshold. Furthermore, the temporal noise (TN) is characterized by the RMS value of the derivative of the S-curve, providing a measure of the sum of both the fluctuations on the injected signal and the fluctuations on the pixel threshold itself. The final threshold value is converted to electrons via the designed value of $C_{\rm test}$.
4.1 Target Working-Point Tuning
To thoroughly assess the impact of reverse bias on sensor performance, it is essential to tune the sensor to target working-points that exhibit a uniform average threshold across various substrate biases. Achieving such consistency poses a challenge due to the complex interplay of multiple parameters influencing the analog front-end. Therefore, the operating conditions—particularly the DAC settings that control the front-end state—must be meticulously adjusted.
This work focuses on the characterization of sectors 0-2. Sector 3 is excluded from the present discussion; its front-end current was increased for radiation tolerance evaluations, leading to significantly higher power consumption. The analog front-end design is identical for sectors 0-2, with variations existing only in their digital logic blocks for pixel functions and in their respective pixel pitches (see Tab. 2).
The M6 MOSFET, biased by VCLIP, clamps excessively high OUT_A signals (see Fig. 1). Given that the amplitude of OUT_A increases significantly with reverse bias due to a reduction in sensor capacitance, VCLIP is set to 1.307 V. This is essentially its maximum value and is chosen to eliminate the clamping effect of M6 on the OUT_A amplitude during characterization. To compensate for the increase of the M8 gate threshold induced by substrate reverse bias, VCASN2 is increased from its nominal 400 mV to 722 mV, thereby restoring M8’s switching threshold to its intended value.
[…]
Table 3: Nominal and Target operating configurations for the analog front-end. The seven Target working-points share constant ITHR, VCLIP, and VCASN2 settings, while VCASN is adjusted to achieve a similar average threshold across different $V_{\rm sub}$ levels.
Settings $V_{\rm sub}$ VCASN/mV ITHR/nA VCLIP/mV VCASN2/mV
Nominal 0 V 400 500 0 400
Target 0 V 400 500 1307 722
-1 V 592
-2 V 732
-3 V 861
-4 V 986
-5 V 1129
-6 V 1265
The threshold distributions for these target working-points are presented in Fig. 5. It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations. This observation is further substantiated by the FPN analysis shown in Fig. 6, where FPN increases with reverse bias and quantitatively reflects the growing differences in threshold across the pixel matrix. The corresponding TN analysis is shown in Fig. 6. An increase in TN is observed under reverse bias conditions. The underlying reasons for this behavior will be discussed in Section 5.2 along with the change of OUT_A signal amplitude.
[…]
Figure 6: (a) Average FPN as a function of reverse bias. (b) Average TN as a function of reverse bias.
4.2 ITHR Scan
To investigate sensor performance across a range of detection thresholds, the ITHR value was systematically scanned starting from each of these target working-points. This procedure yielded a series of operating conditions—7 distinct $V_{\rm sub}$ levels, each with 11 different ITHR settings—resulting in a set of working-points with varying thresholds.
Figure 7 illustrates the impact of the ITHR value on the mean threshold, FPN and TN, respectively. Both FPN and TN generally exhibit an increasing trend with higher threshold values.
The linear relationship between the mean threshold and the ITHR setting, depicted in Fig. 7, is consistently observed across all tested reverse bias voltages. This linearity not only confirms the consistency of the threshold tuning but also underscores that ITHR is a suitable parameter for threshold control under different reverse biases.
Figure 7: ITHR scan results. (a) Threshold, (b) FPN (fixed pattern noise), and (c) TN (temporal noise) versus ITHR value for working-points of distinct reverse biases.
5 Experimental Results and Discussion
Reducing sensor capacitance and expanding the depletion zone are the most intrinsic and direct consequences of applying a reverse bias on the substrate. After adjusting thresholds under different reverse biases and determining a set of working-points by ITHR value scan described in Sec. 4.2, a thorough sensor characterization regarding the impact of the substrate reverse bias voltage was conducted, including leakage current, analog response, fake-hit rate and charge collection efficiency.
5.1 Leakage Current
In the absence of ionizing radiation, thermally generated free charge carriers are collected by the diode, resulting in what is known as leakage current or dark current. Leakage current from the pixel area impacts the sensor’s electrical performance by inducing shot noise [18].
[…]
5.2 Input Capacitance
The effective pixel input capacitance of the sensor, $C_{\rm effective}$, can be determined by analyzing the analog response of the sensor under different reverse bias voltages. The analog output amplitude (denoted as OUT_A in Fig. 1) is given by:
$A_{\rm{OUT\_A}}=\frac{C_{\rm test}\cdot V_{\rm pulse}}{C_{\rm effective}}\cdot f_{\rm gain}\cdot f_{\rm sf}~{},$ (1)
[…]
Two independent and complementary methods were employed in this study to measure $C_{\rm effective}$ via the relation:
$C_{\rm effective}=C_{\rm test}\cdot\frac{V_{\rm pulse}}{A_{\rm{OUT\_A}}}\cdot f_{\rm gain}\cdot f_{\rm sf}~{}.$ (2)
  • 1.
    Method 1: In this approach, the pulse voltage was fixed at $V_{\rm pulse}=160{\rm mV}$ while the analog response amplitude was recorded via an oscilloscope. The measurement was repeated to extract a statistically robust value for the amplitude. The measured amplitude is then substituted into Eq. (2), the term $A_{\rm{OUT\_A}}$, to obtain $C_{\rm effective}$.
  • 2.
    Method 2: In this procedure, the analog response amplitude, $A_{\rm{OUT\_A}}$, was maintained at a constant value by dynamically adjusting $V_{\rm pulse}$. The corresponding waveform measurement was repeated, and the adjusted $V_{\rm pulse}$ value (required to hold the amplitude constant) was then used in Eq. (2) to compute $C_{\rm effective}$.
Figure 9: Effective capacitance measured as a function of the reverse bias voltage using two methods. (a)Method 1: $V_{\rm pulse}$ is fixed at 160 mV, and the analog response amplitude is measured over 200 repetitions. (b) and (c)Method 2: The analog amplitude is fixed at 49 mV, 78 mV, respectively, by adjusting $V_{\rm pulse}$; measurements are performed at two different current threshold settings (500 pA and 650 pA, respectively). For both methods, the calculation of $C_{\rm effective}$ assumes $f_{\rm gain}=20$ and $f_{\rm sf}=0.7$.
The two methods provide complementary insights. While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage. To examine how $f_{\rm gain}$ varies with amplitude, Method 2 was repeated at two different output levels (49 mV and 78 mV), shown in Fig. 9 and Fig. 9. The results show consistent trends with only tiny differences between the two amplitudes, confirming the reliability of the approach.
Results for these two methods after calculation are shown in Fig. 9. The input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{}{\rm V}$ to $V_{\rm sub}=-2~{}{\rm V}$, corresponding to the rapid expansion of the depletion region. This reduction is very important as smaller input capacitance increases the charge-voltage conversion coefficient at the input of the front-end circuit for a given charge. To keep the input-referred threshold constant, the VCASN was tuned to lower the OUT_A voltage baseline as the OUT_A amplitude increased with reverse bias voltage. This may be related to the increase of FPN and TN for reverse bias from -1 V to -5 V in Fig. 6 6 respectively, as the same effect had been observed in a VCASN scan shown in Ref. [11] (Fig. 13).
5.3 Fake-hit Rate Test
Fake-hit, i.e., spurious signals that do not correspond to any external particle interactions, are a critical parameter in pixel detector performance. These false signals often arise from electronic noise, crosstalk, or other sensor imperfections, leading to erroneous track reconstruction. Therefore, assessing and reducing the fake-hit rate is crucial for ensuring the reliability of tracking detectors.
To compare two distinguished readout sequence modes, the fake-hit rate for each sector $i$ in both rolling shutter (RS) and global shutter (GS) modes is calculated using a unified formula:
$R_{i}^{\text{mode}}=\frac{N_{i}^{\text{mode}}}{N_{\text{frames}}^{\text{mode}}\cdot N_{\text{pixels}}},$ (3)
where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector. In the rolling shutter mode (RS), 100,000 frames are collected ($N_{\text{frames}}^{\text{RS}}=100,000$) in the absence of external sources. In the global shutter mode (GS), 1,000 frames are collected ($N_{\text{frames}}^{\text{GS}}=1,000$) with an acquisition window of 98.3 $\upmu\rm s$, matching the rolling shutter readout time, to enable a comparable measurement.
The sensitivity limit is defined as the lowest measurable fake-hit in the measurement for these two modes and is selected as $\rm 1\times 10^{-10}$ for the rolling shutter and $\rm 1\times 10^{-8}$ for the global shutter mode.
As the fake-hit rate is expected to be strongly influenced by the threshold, a threshold scan under different reverse biases was performed. The results are shown in Fig. 10. In Fig. 10, up to 10 pixels (accounting for 0.04%) in each sector with the highest fake-hit counts are identified as hot pixels that contribute most to the false hits [20, 21]. These hot pixels are masked. In comparison, values presented in Fig. 10 are obtained without masking any pixels.
According to these results, the difference between the two modes is insignificant above the GS sensitivity limit. Reverse bias substantially reduces the fake-hit rate compared with the unbiased state, with $V_{\rm sub}=-3~{}{\rm V},-4~{}{\rm V}$ yielding the lowest values.
Figure 10: Fake-hit Rate measured with RS (rolling shutter) and GS (global shutter) modes versus threshold under varied reverse biases with all pixels included in Fig. 10 and with hot pixels masked in Fig. 10. In Fig. 10, up to 10 pixels (accounting for 0.04%) for each sector, which have the largest counts of fake-hit, are considered hot pixels and masked. Sensitivity limits are indicated as dashed horizontal lines in every pad.
5.4 Radioactive Source Test
Two radioactive sources, ${}^{90}\rm Sr$ and ${}^{55}\rm Fe$, were employed to assess the sensor’s response to external ionizing radiation. ${}^{90}\rm Sr$, a $\beta$ emitter, generates minimum ionizing particles that traverse the sensor and deposit charge over an extended region. This distributed ionization is particularly sensitive to changes in the sensor’s depletion characteristics and charge-sharing behavior under reverse bias, thereby providing insights into the sensor’s charge collection efficiency. In contrast, ${}^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition. This characteristic makes ${}^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization.
5.4.1 Test with ${}^{90}\rm Sr$ Source
A ${}^{90}\rm Sr$ source was placed 2 mm above the sensor to verify JadePix-3’s response to beta rays. All data were collected with a 100,000-frame rolling shutter scan, ensuring equivalent collection time for every tested working-points. Approximately 30,000 clusters were recorded in sector 0,1,2 over the course of 50 seconds, for which the hit map is shown in Fig. 11. Fig. 12 displays the data analysis result. Due to varying solid angles relative to the fixed source, each sector inherently accumulates a different number of hits. However, because each working-point was recorded with the same number of frames (and thus identical detection time), and radioactive source was kept stationary throughout the measurements, the accumulated cluster count within a given sector is directly proportional to its detection efficiency. Therefore, for each sector, the relative detection efficiency at a given bias voltage was obtained by normalizing its accumulated cluster count by the maximum cluster count recorded for that same sector across all bias settings.
Figure 11: Hit map of sector 0,1,2 from the ${}^{90}\rm Sr$ radioactive source test. Roughly 30,000 events were recorded under $\beta$-ray irradiation from the backside during a time duration of 50 s, with ITHR value fixed at 500 pA and $V_{\rm sub}$ at 0 V.
Figure 12: ${}^{90}\rm Sr$ radioactive source response result. (a) Average cluster size versus the threshold. (b)-(d) Relative detection efficiency versus the threshold for Sector 0, 1, 2, respectively. The relative detection efficiency is proportional to the accumulated number of clusters detected and normalized by the maximum number of clusters measured in each sector.
Fig. 12 demonstrates a clear reduction in cluster size with the increase of the bias voltage, e.g., from 16-17 to 8-9 at 100 $e^{-}$ threshold from 0 V to -6 V. This means more charge from particle interactions is collected within the seed pixel, reducing the spread to neighboring pixels. Additionally, the stronger electric field lowers lateral diffusion, causing the charge to be collected more locally. As a result, fewer pixels surrounding the seed pixel exceed the detection threshold, leading to smaller average cluster size. The slope of cluster size as a function of the threshold also decreases significantly in Fig. 12. This phenomenon demonstrates that the charge collection ability of the sensor is less affected by the threshold under biased conditions, which further illustrates that the charge collection is more sufficient and robust.
Fig. 12-12 demonstrate an enhancement in detection efficiency as reverse bias increases from 0 V to -5 V. This behavior can be attributed to the depletion zone expansion. However, the measured results also exhibit an evident efficiency drop as the bias is further increased from –5 V to –6 V. One possible reason for this decline is that under excessive reverse bias, the characteristics of the MOSFETs may undergo significant changes due to the body effect, leading to a reduction in signal amplification performance. A similar effect under high bulk bias was also observed in Ref. [22], where an nMOS device biased at –6 V exhibited a 260 mV threshold voltage shift and a 40% drop in drain current. This suggests that the analog front-end performance of JadePix-3 may be compromised under such high bias conditions, highlighting the importance of carefully selecting the operational bias range.
5.4.2 Test with ${}^{55}\rm Fe$ Source
A ${}^{55}\rm Fe$ source was positioned 8 mm above the sensor to verify JadePix-3’s response to X-ray. The data collection and processing for ${}^{55}\rm Fe$ were identical to ${}^{90}\rm Sr$, using a 100,000-frame rolling shutter scan, recording approximately 30,000 clusters over 50 seconds, ensuring a consistent collection time window.
Figure 13: ${}^{55}\rm Fe$ radioactive source response result. (a) Average cluster size versus the threshold. (b)-(d) Relative detection efficiency versus the threshold of Sector 0, 1, 2, respectively. The relative detection efficiency is obtained in the same method as for the ${}^{90}\rm Sr$ source test as shown in Fig. 12.
As depicted in Fig. 13, the slope of the cluster size versus threshold curve decreases significantly when $V_{\rm sub}$ increases, providing similar results to the ${}^{90}\rm Sr$ source test, which further illustrates the charge collection is more sufficient under the biased condition. Fig. 13-13 shows improvement in detection efficiency as substrate reverse bias increases.
Comparing Fig. 13-13 internally, sector 1 exhibits a smaller efficiency drop with increasing the threshold (approximately 10%) compared to sectors 0 (20%) and 2 (25%). A similar behavior is also observed for the ${}^{90}\rm Sr$ source test, as shown in Fig. 121212. A plausible mechanism involves the difference of the OUT_D trigger design: sector 1 uses a D-flipflop, while both sector 0 and 2 use SR-flipflops (see Tab. 2). D-flipflop is edge-triggered. While the SR-flipflop is level-triggered or latched. The discrepancy may indicate that the electrical test pulses are different from the ionization signal because the thresholds are calibrated by the electrical pulse while they showed different results on source measurements.
5.5 Infrared Laser Test
Following the configuration of the laser test platform outlined in Section 3, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam. The laser operates at a wavelength of 1064 nm. The laser’s pulse energy is adjustable between 0 and 60 pJ, modulated via an attenuation factor ranging from 0% tune (maximum pulse energy) to 100% tune (minimum pulse energy), controlled through the Laser Power Tune function of the system. For this laser measurement, the selected pixel is located in sector 2, which has the smallest pixel pitch among the sectors (see Tab. 2).
5.5.1 Beam Position Adjustment
To ensure optimal focusing and achieve the minimum laser spot size, which can be tuned to 3.4 $\rm{\upmu m}$, significantly smaller than the pixel size [11], at the depth of the chip’s sensitive layer, a scan was performed along the Z-direction (perpendicular to the sensor plane) to locate the beam waist. This scan involves measuring the number of pixel hits as a function of the laser’s Z-position, which indicates the vertical distance to the sensor surface, at two attenuation levels: 5% and 50%. As shown in Fig. 14, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm. This position was subsequently fixed for all subsequent measurements to ensure optimal focus and minimal spot size on the chip.
Figure 14: Number of pixel hits as a function of the laser position Z. The upper and lower curves are for the laser power attenuation at 5% and 50%, respectively. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frame numbers = 1,000.
To approach the center of a single pixel, a two-dimensional XY-scan along the sensor surface was performed, as illustrated in Fig. 15. The cluster size during this scan was restricted to either 0 or 1 under a 75% laser tune setting, ensuring precise delineation of the pixel boundaries. The experimental conditions included a laser frequency of 25 kHz, a threshold of 215 $e^{-}$ for the sensor, a substrate bias voltage $V_{\rm sub}$ of 0 V, and a total of 10,000 Rolling Shutter frames. This scan confirmed the exact center position of a single pixel in both the X- and Y-directions.
Figure 15: XY-scan results from a laser source. The plot on the left shows the hit map, the number of laser pulse events recorded at a specific (X, Y) position. The upper right panel displays the hit profile along the line at Y = 9.774 mm, while the lower right panel shows the hit profile along the line at X = 20.263 mm. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frames = 10,000.
5.5.2 In-Pixel Detection Efficiency
Building on previous beam test results [13], we further probed the in-pixel detection performance by using the attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency in response to external excitation.
Based on the XY-scan results as shown in Fig. 15, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels. These four positions, annotated in Fig. 15, were representative of distinct in-pixel locations and were selected to investigate spatial variations in detection efficiency.
As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s detection efficiency in response to it to mimic the pathways of charged particles. An RS scan was performed over 100,000 frames (40 seconds) under a 25 Hz laser operation for each reverse bias value across the four in-pixel positions. Given that 1000 laser pulse triggers are expected during this period, the detection efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses. Results shown in Fig. 16 demonstrate that applying a reverse bias from 0 to –2 V had already enhanced the detection efficiency at in-pixel positions A, B, and C to nearly 100%, underscoring the performance improvement enabled by the reverse bias. These observations are in agreement with beam tests of other MAPS sensors [23], providing an imperative reference to further beam test results.
Figure 16: Detection efficiency in response to laser beam as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 15.
Furthermore, a laser with tuning method was developed and performed in order to quantitatively characterize the sensor’s charge collection capability. The most straightforward way would be to measure the charge deposit in the sensor under external excitation, however, JadePix-3 employs a binary digital readout, and it lacks the analog output of the collected charge. Therefore, an external probe was required whose intensity can be continuously varied to infer the pixel’s charge collection sensitivity. By using the laser’s tunable attenuation as a proxy for the amount of injected charge, the sensor’s binary readout can be translated into a quantitative efficiency index in a method described as follows. The laser frequency was set to 25 Hz, and the chip was operated with 20,000 Rolling Shutter frames, equivalent to an 8-second acquisition window. Under these conditions, if every laser pulse generates enough charge that exceeds the detection threshold, 200 hits would be recorded by the pixel. The laser pulse energy was then attenuated by incrementally increasing the tune value until the number of recorded hits dropped to around 100333The actual experimental values are not exactly 100 and range between 90 and 110 due to unavoidable fluctuation.. The tune value at this point, designated as the Detection Sensitivity Index (DSI), was recorded for each of the four marked positions. This DSI is defined as the laser tune value at which the number of detected hits drops to approximately 100 out of 200 possible triggers; a higher DSI indicates that the pixel remains responsive even at lower laser intensities, thus signifying more effective charge collection efficiency and thus improved detection performance.
Figure 17: Detection Sensitivity Index (DSI) as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 15.
The variation of the DSI with reverse bias $V_{\rm sub}$ was measured and is depicted in Fig. 17. The resulting curves reveal distinct trends across the four in-pixel positions, providing more intricate information about the in-pixel variance. Position A, the pixel center, consistently exhibited the highest DSI, indicating peak detection efficiency, whereas Position D showed the lowest DSI, reflecting diminished sensitivity at the pixel boundary. Position B (long-edge center) and C (short-edge center) displayed intermediate DSI, with Position B substantially outperforming Position C. This aligns with findings from the JadePix-3 beam test measurement, which revealed notably higher detection efficiency at the long-edge center (Position B) compared to the short-edge center (Position C), where efficiency was significantly lower and closely resembled that of the pixel corner (Position D). Notably, the DSI increased with increasing reverse bias across all positions, suggesting enhanced charge collection efficiency at higher $V_{\rm sub}$ due to the expanding depletion volume.
5.6 Optimal Reverse Bias Setting
A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that a reverse bias of approximately –5 V offers the best overall trade-off. This optimal working-point not only improves detection efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
6 Conclusions and Outlook
To further characterize the performance of JadePix-3, a prototype MAPS developed for the CEPC vertex detector, detailed laboratory-based measurements were performed, in particular with substrate reverse bias voltages applied. A comprehensive test workflow was developed and implemented for the systematic characterization of JadePix-3. Experimental results show that the JadePix-3 sensor can work stably when a substrate reverse bias is applied. Moreover, the reverse bias has a considerable influence on the performance of the sensor, typically being beneficial. Our results reveal that applying reverse bias substantially enhances sensor performance, reducing input capacitance, enlarging the depletion region, and improving charge collection efficiency, while also lowering fake-hit rates. Optimal reverse bias setting is also determined within this integrated test procedure. This test methodology not only confirms JadePix-3’s potential for further performance improvement but also establishes a robust methodology for the evaluation and optimization of similar MAPS devices in future particle tracking applications.
This result acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test. Moreover, this work will serve as the architectural and procedural foundation for the tuning, and optimization of the next iteration, the JadePix-5 sensor, thereby driving forward MAPS development to meet CEPC’s requirements.
Acknowledgement
This work was supported in part by the National Key Research and Development Program of China under Grant 2022YFA1602203 and Grant 2023YFA1606300, the Fundamental Research Funds for the Central Universities of China under Grant WK2360000014, and the STCF Key Technology Research and Development Project. We thank the Hefei Comprehensive National Science Center for its strong support on the promotion and development of the STCF project.

### v2

@@
-Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector
+Further Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector
 Abstract
 The Circular Electron-Positron Collider (CEPC), a proposed next-generation $e^{+}e^{-}$ collider to enable high-precision studies of the Higgs boson and potential new physics, imposes rigorous demands on detector technologies, particularly the vertex detector.
-JadePix-3, a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector, targets a spatial resolution of 3 $\rm{\upmu m}$ to meet the key requirement.
+JadePix-3 is a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector.
 This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and detection efficiency.
 Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate.
-Simultaneously, an extensive methodology for MAPS characterization is established, which determined the optimal operational working conditions for the JadePix-3 sensor.
-Furthermore, these findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
+These findings confirm the sensor’s potential for high-precision particle tracking and vertexing at the CEPC while offering valuable references for future iterational R&D of the JadePix series.
 keywords:
@@
 […]
+{highlights}
+Comprehensive laboratory characterization of the JadePix-3 CMOS pixel prototype for the vertex detector of the Circular Electron-Positron Collider
+Measurements performed under various reverse bias voltages from 0 to –6 V with dedicated working point optimization at each bias voltage
+Substrate reverse bias significantly improves the sensor performance such as the charge collection efficiency and the fake-hit rate
 1 Introduction
-The discovery of the Higgs boson has spurred interest in circular $e^{+}e^{-}$ colliders as premier facilities for high-precision studies of the Standard Model Higgs boson and other fundamental particles.
-Planned designs aim to cover a wide range of center-of-mass energies—from the Z pole (91 GeV) and WW threshold (161 GeV) to the Higgs factory region (240 GeV) and the $t\bar{t}$ threshold (340–365 GeV)—using an $e^{+}e^{-}$ accelerator with a circumference of approximately 100 km.
-In addition, the same tunnel could later accommodate a hadron collider with a center-of-mass energy nearly 100 TeV.
-The Circular Electron-Positron Collider (CEPC) [1, 2] has been proposed to pursue this road map, envisioned to investigate the Higgs boson further and reveal new physics potentially.
-CEPC’s ambitious physics objectives impose stringent performance requirements on its detectors—particularly the vertex detector.
-As outlined in Table 1, the CEPC vertex detector must achieve exceptional spatial resolution, maintain a low material budget, minimize power consumption, handle high hit densities, and exhibit robust radiation tolerance to cope with the beam-induced background near the interaction point.
-In this context, only Monolithic Active Pixel Sensors (MAPS) offer a promising solution to meet these demanding criteria [3, 4].
-Therefore, considerable R&D efforts have been undertaken to develop advanced pixel sensor prototypes with monolithic technology for the CEPC vertex detector, such as the TaichuPix series [5, 6, 7] and the JadePix series.
-The JadePix series is specifically developed to meet the CEPC vertex detector’s demanding criteria, offering high position resolution along with enhanced efficiency.
-Building on the progressive improvements of its predecessors—JadePix-1, which was to verify the sensor fabrication and optimize the pixel architecture [8, 9], and JadePix-2, which featured an improved analog front‐end design [10]—JadePix-3 represents the latest development and aims for a position resolution of 3 $\rm{\upmu m}$ [11].
-The next part presents the detailed design specifications of JadePix-3.
-Table 1: Demands for the CEPC vertex detector [1].
+Following the discovery of the Higgs boson, circular $e^{+}e^{-}$ colliders have garnered significant interest as premier facilities for high-precision studies of the Standard Model and searches for new physics.
+The Circular Electron-Positron Collider (CEPC) [1, 2] has been proposed to advance this frontier, with an ambitious physics program centered on detailed investigations of the Higgs boson.
+The ambitious physics objectives of the CEPC impose stringent performance requirements on its detector systems, particularly on the vertex detector located nearest to the interaction point.
+As outlined in Table 1, the vertex detector must satisfy a challenging combination of specifications: exceptional spatial resolution, a minimal material budget, low power consumption, high hit-rate capability, and robust radiation tolerance against beam-induced backgrounds.
+Meeting these multifaceted demands is a significant technological challenge.
+In this context, Monolithic Active Pixel Sensors (MAPS) stand out as a uniquely promising technology capable of fulfilling all these criteria [3, 4].
+Consequently, considerable research and development efforts are underway to advance monolithic sensor prototypes for the CEPC.
+The JadePix sensor series, developed specifically for this purpose, is a direct result of these efforts.
+This paper focuses on the JadePix-3 prototype, the design specifications of which are briefly summarized below, with further details available in Ref.
+[5].
+Table 1: Requirements for the CEPC vertex detector [1].
 Single point resolution $\rm{\upmu m}$
@@
 Radiation tolerance TID, 3.4 Mrad/year
-NIEL, $\rm 6.2\times 10^{12}n_{eq}/(cm^{2}year)$
-Figure 1: (a) Layout of the entire sensor mask; (b) Circuit topology for the analog front-end [11].
+NIEL, $\rm 6.2\times 10^{12}~\text{1~MeV}~n_{eq}/(cm^{2}\cdot year)$
 1.1 Overview of JadePix-3 design
-The JadePix-3 is designed for a high spatial resolution of 3 $\rm{\upmu m}$.
-With a pixel matrix consisting of 512 rows and 192 columns, the sensor covers a sensitive area of 8.19 $\times$ 4.85 $\rm{mm^{2}}$ as shown in Fig.
-1.
+The JadePix-3 sensor, developed with TowerJazz 180 nm CMOS imaging technology, covers a pixel matrix consisting of 512 rows and 192 columns.
 The sensor is divided into four distinguishable sectors in the column direction.
-As presented in Tab.
-2, while the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3.
-The pixel arrangement in each sector is mirrored, with shared vertical metal lines for biasing and control signals.
-Fig.
-1 demonstrates JadePix-3’s front-end circuit design.
-As described in Ref.
-[11] in detailed, the initial amplification stage utilizes a cascode amplifier, which provides high gain while minimizing noise, with its output referred to as OUT_A.
-Subsequently, the amplified signal is fed into a discriminator stage, where it is compared against a predefined threshold.
-The resulting digital output, designated as OUT_D, enables efficient hit detection.
+While the row pitch is uniformly set at 16 $\rm{\upmu m}$ across all four sectors, the column pitch varies between 23.11 $\rm{\upmu m}$ in sector 2 and 26 $\rm{\upmu m}$ in sectors 0, 1, and 3.
+The variation in the column pitch arises from the implementation of different digital logic blocks within the pixels.
+These blocks were designed to evaluate the performance of R-S flipflops versus D flipflops, as well as the inclusion or exclusion of a MASK configuration register, requiring adjustments to the pixel layout.
+All these differences are summarized in Tab.
+2.
 Figure 1: Schematic diagram of the cross-section of the sensor when a reverse bias is applied to the P-substrate via a P-well.
@@
 Sector pixel layout digital logic block Remarks
-0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 SR flipflop
+0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 R-S flipflop
 1 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V1 D flipflop, w/o MASK
-2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 SR flipflop, w/o MASK
+2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 R-S flipflop, w/o MASK
 3 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 Increased current for the front-end
-The chip employs both rolling shutter (RS) and global shutter (GS) readout sequence modes.
-In the conventional rolling shutter mode, one row is activated for reading at a time [12].
-This process requires 192 $\rm ns$ to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$.
-While in global shutter, all 512 rows are frozen for a period of time during readout.
-The JadePix-3’s sensor design is shown in Fig. 2. A reverse bias $V_{\rm sub}$ can be applied on P-well and P-substrate, expanding the depletion region around the N-well electrode.
+In the JadePix-3 sensor design, each pixel contains a P-well that allows for substrate biasing through a surface contact, as shown schematically in Fig.
+1.
+This P-well is electrically contiguous with the underlying deep P-well and the P-type epitaxial layer.
+This configuration ensures that the entire P-type substrate is held at the applied potential, $V_{\rm sub}$.
+The pads connected to these P-wells are designed without ESD protection to allow for the application of a wide range of substrate bias voltages.
+When a reverse bias $V_{\rm sub}$ is applied on P-well and P-substrate, the depletion region around the N-well electrode expands.
 1.2 Characterization Motivation
@@
 However, the full performance potential of the JadePix-3 sensor remains incompletely characterized, particularly concerning the effects of reverse bias.
+Figure 2: Circuit topology for the JadePix-3 analog front-end [5].
 The performance of MAPS is significantly influenced by the bias voltage applied to the substrate.
-This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [14, 15].
-On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry tends to increase due to the body effect, consequently altering the circuit’s operating status.
+This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [7].
+On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry (shown in Fig.
+2) tends to increase due to the body effect, consequently altering the circuit’s operating status.
 Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential.
@@
 Consequently, this work provides a thorough study of reverse bias effects, which is imperative for optimizing sensor design and ensuring its suitability for high-precision particle tracking applications.
-This work also introduces an integrated and systematic test methodology (detailed in Section 2), which was established to guide the subsequent comprehensive, lab-based characterization of MAPS behavior under varying reverse bias conditions.
-While most studies hold the substrate bias fixed or explore a narrow range [16, 17], our approach systematically sweeps the reverse bias from 0 V to –6 V under identical threshold settings.
+This work also introduces an integrated and systematic test methodology (detailed in Section 2.2), which was applied to guide the subsequent lab-based characterization of MAPS behavior under varying reverse bias conditions.
+Our approach systematically sweeps the reverse bias from 0 V to –6 V under identical threshold settings.
 This methodology establishes comparable operating conditions across different bias voltages, thereby facilitating an unambiguous assessment of performance changes predominantly attributable to the bias variation itself.
-2 Workflow for Comprehensive Characterization
-This work employs a structured, integrated methodology (see Fig. 3):
-  • 1.
-    Phase 1: Measurement Setup, which is introduced in Section 3, focuses on the preparation stage and establishes a stable experimental setup that minimizes mechanical and environmental perturbations, thereby ensuring reliable test conditions.
-  • 2.
-    Phase 2: Working-Point Tuning is introduced in Section 4. This phase establishes a series of optimal operating points through a two-stage procedure.
-    1. (a)
-      Target threshold: Target working-points are defined as distinct operating configurations, one for each applied bias level, that all share a consistent average threshold.
-This is achieved by applying reverse bias and meticulously tuning DAC parameters to meet the predefined target threshold.
-    2. (b)
-      ITHR scan: Subsequently, an ITHR scan is performed from each of these target working-points to generate an expanded set of conditions, offering a range of distinct threshold levels for each bias.
-    This tuning not only confirms the feasibility of operating under varying reverse biases but also lays the groundwork for all subsequent measurements.
+2 Experimental Methods
+2.1 Measurement Setup
+The experimental setup, which is based on the one described in Ref.
+[5], is shown in Fig.
 3.
-    In Phase 3: Comprehensive Performance Evaluation, which is illustrated in Section 5, three sequentially interlinked investigations were carefully designed and performed:
-    1. (a)
-      Sensor Electrical Properties Characterization: Detailed measurements of leakage current and input capacitance—both regarded as critical sensor performance metrics—were conducted to understand the intrinsic electrical behavior of the sensor under bias.
-    2. (b)
-      Noise Performance Analysis: Fake-hit rate is measured to assess noise characteristics, since maintaining a low fake-hit rate is critical for ensuring the effectiveness of hit data for the subsequent measurements.
-    3. (c)
-      Detection Response Evaluation: The sensor’s response is further assessed by exposing it to external sources, including radioactive sources and an infrared laser.
-The laser tests, in particular, provide high-resolution mapping of in-pixel charge collection efficiency and serve as a proxy for evaluating sensor performance in response to external stimuli.
-Figure 3: Detailed workflow to comprehensively characterize the MAPS sensor.
-This integrated workflow not only delivers a robust, reproducible characterization of JadePix-3 in the laboratory—fully capturing how reverse bias shapes every aspect of its response—but also provides a transferable framework for future MAPS R&D.
-3 Experimental Setup
-The JadePix-3 chip has been wire-bonded to an adaptor board and then connected to an external FPGA-based DAQ system (see Fig.
-4).
+The JadePix-3 chip has been wire-bonded to a carrier board and then connected to an external FPGA-based DAQ system (see Fig.
+3).
 The FPGA board configures the chip parameters and transfers data to the PC via Ethernet.
@@
 A source meter is used to apply reverse bias ranging from 0 to -6 V to the substrate.
-Figure 4: (a) Schematic view of the test system; (b) Setup of the infrared laser test system. All devices are fixed or placed on an optical table.
-To proceed laser experiment under the demand of high spatial resolution, a laser test system featuring three high-precision motor controller is deployed (see Fig.
-4).
+The upper limit of -6 V was chosen as a conservative value to ensure a stable operation, based on breakdown voltages reported for other sensors fabricated in the same 180 nm process [7].
+Figure 3: (a) Schematic view of the test system; (b) Setup of the infrared laser test system.
+There is a window in the PCB to allow radiative source or laser beam incident to the JadePix-3 sensor.
+All devices are fixed or placed on an optical table.
+To conduct laser experiment under the demand of high spatial resolution, a laser test system featuring three high-precision motor controllers is deployed (see Fig.
+3).
 The spatial position in three orthogonal directions of the laser source is regulated precisely by these three motors, each has a repeatable position resolution of 1 $\rm{\upmu m}$ and a minimum step size of 0.1 $\rm{\upmu m}$.
-Details about the experiment parameters and test method will be discussed thoroughly in Section 5.5.
-4 Working-point Selection
-This section details the methodology for selecting operational working-points of the JadePix-3 sensor.
+Details about the experiment parameters and test method will be discussed in Section 3.5.
+2.2 Working-point Selection
+This part details the methodology for selecting operational working-points of the JadePix-3 sensor.
 The primary objective is to establish uniform thresholds that enable the comparison of the sensor’s performance across varying substrate reverse biases.
 Central to this process is the accurate measurement of pixel thresholds, achieved using the S-curve method.
-This methodology involves giving each pixel a certain number of electrical pulses and counting the hits while increasing the injected charge, which is regulated as $V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor that has a designed value of $0.18~{}{\rm fF}$ marked in Fig.
-1, and $V_{\rm pulse}$ is the external voltage for test-pulse injection.
-As the injected charge approaches the pixel threshold, the hit count rises along an S-shaped curve, typically modeled by fitting an error function.
-The pixel threshold is then defined as the injected charge level at which this fitted error function reaches half of its maximum value.
-Threshold variations between pixels are quantified by the fixed pattern noise (FPN), defined as the standard deviation of the mean threshold.
-Furthermore, the temporal noise (TN) is characterized by the RMS value of the derivative of the S-curve, providing a measure of the sum of both the fluctuations on the injected signal and the fluctuations on the pixel threshold itself.
-The final threshold value is converted to electrons via the designed value of $C_{\rm test}$.
-4.1 Target Working-Point Tuning
+The operational parameters of the pixels—threshold, fixed pattern noise (FPN), and temporal noise (TN)—were characterized using the standard “S-curve” method [8].
+This technique involves injecting a series of electrical pulses of increasing charge ($Q_{\rm injected}$) into each pixel and measuring the hit response.
+The injected charge is regulated by an external voltage pulse, $V_{\rm pulse}$, according to the relation $Q_{\rm injected}=V_{\rm pulse}\cdot C_{\rm test}$, where $C_{\rm test}$ is the injection capacitor with a designed value of 0.18 fF.
+The resulting S-shaped curve for each pixel is then analyzed to extract its specific threshold and noise characteristics.
+The final threshold value is converted to electrons via the value of $C_{\rm test}$.
+2.2.1 Target Working-Point Tuning
 To thoroughly assess the impact of reverse bias on sensor performance, it is essential to tune the sensor to target working-points that exhibit a uniform average threshold across various substrate biases.
 Achieving such consistency poses a challenge due to the complex interplay of multiple parameters influencing the analog front-end.
-Therefore, the operating conditions—particularly the DAC settings that control the front-end state—must be meticulously adjusted.
+Therefore, the operating conditions—particularly the DAC settings that control the front-end bias condition—must be meticulously adjusted.
 This work focuses on the characterization of sectors 0-2.
-Sector 3 is excluded from the present discussion; its front-end current was increased for radiation tolerance evaluations, leading to significantly higher power consumption.
+Sectors 0-2 share the common DAC configuration.
 The analog front-end design is identical for sectors 0-2, with variations existing only in their digital logic blocks for pixel functions and in their respective pixel pitches (see Tab.
 2).
+Sector 3 is excluded from the present discussion; its front-end current was increased for radiation tolerance evaluations, leading to significantly higher power consumption.
 The M6 MOSFET, biased by VCLIP, clamps excessively high OUT_A signals (see Fig.
-1).
-Given that the amplitude of OUT_A increases significantly with reverse bias due to a reduction in sensor capacitance, VCLIP is set to 1.307 V.
-This is essentially its maximum value and is chosen to eliminate the clamping effect of M6 on the OUT_A amplitude during characterization.
-To compensate for the increase of the M8 gate threshold induced by substrate reverse bias, VCASN2 is increased from its nominal 400 mV to 722 mV, thereby restoring M8’s switching threshold to its intended value.
+2).
+Given that the amplitude of OUT_A increases significantly with reverse bias due to a reduction in the sensor capacitance, VCLIP is set to 1.307 V.
+This is its maximum value and is chosen to eliminate the clamping effect of M6 on the OUT_A amplitude during characterization.
+To compensate for the increase of the M8 gate threshold induced by substrate reverse bias, VCASN2 is increased from its nominal 400 mV to 722 mV, thereby maintaining a sufficient gate voltage for the cascode structure of M8 and M9.
 […]
@@
 The seven Target working-points share constant ITHR, VCLIP, and VCASN2 settings, while VCASN is adjusted to achieve a similar average threshold across different $V_{\rm sub}$ levels.
-Settings $V_{\rm sub}$ VCASN/mV ITHR/nA VCLIP/mV VCASN2/mV
+Settings $V_{\rm sub}$ VCASN/mV ITHR/pA VCLIP/mV VCASN2/mV
 Nominal 0 V 400 500 0 400
-Target 0 V 400 500 1307 722
+Target
+0 V 400 500 1307 722
 -1 V 592
@@
 The threshold distributions for these target working-points are presented in Fig.
+4.
+It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations.
+This broadening of the threshold distributions is quantified in Fig.
+5, which plots the FPN as a function of bias voltage.
+The FPN is observed to increase with reverse bias, reflecting the growing pixel-to-pixel variations.
+The corresponding TN analysis is shown in Fig.
 5.
-It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations.
-This observation is further substantiated by the FPN analysis shown in Fig.
-6, where FPN increases with reverse bias and quantitatively reflects the growing differences in threshold across the pixel matrix.
-The corresponding TN analysis is shown in Fig.
-6.
 An increase in TN is observed under reverse bias conditions.
-The underlying reasons for this behavior will be discussed in Section 5.2 along with the change of OUT_A signal amplitude.
+The underlying reasons for this behavior will be discussed in Section 3.2 along with the change of the OUT_A signal amplitude.
 […]
 Figure 5: (a) Average FPN as a function of reverse bias. (b) Average TN as a function of reverse bias.
-4.2 ITHR Scan
-To investigate sensor performance across a range of detection thresholds, the ITHR value was systematically scanned starting from each of these target working-points.
+2.2.2 ITHR Scan
+To investigate the sensor performance across a range of detection thresholds, the ITHR value was systematically scanned starting from each of these target working-points.
 This procedure yielded a series of operating conditions—7 distinct $V_{\rm sub}$ levels, each with 11 different ITHR settings—resulting in a set of working-points with varying thresholds.
-Figure 7 illustrates the impact of the ITHR value on the mean threshold, FPN and TN, respectively. Both FPN and TN generally exhibit an increasing trend with higher threshold values.
+Fig. 6 illustrates the impact of the ITHR value on the mean threshold, FPN and TN, respectively. Both FPN and TN generally exhibit an increasing trend with higher threshold values.
 The linear relationship between the mean threshold and the ITHR setting, depicted in Fig.
-7, is consistently observed across all tested reverse bias voltages.
+6, is consistently observed across all tested reverse bias voltages.
 This linearity not only confirms the consistency of the threshold tuning but also underscores that ITHR is a suitable parameter for threshold control under different reverse biases.
@@
 3 Experimental Results and Discussion
-Reducing sensor capacitance and expanding the depletion zone are the most intrinsic and direct consequences of applying a reverse bias on the substrate.
-After adjusting thresholds under different reverse biases and determining a set of working-points by ITHR value scan described in Sec.
-4.2, a thorough sensor characterization regarding the impact of the substrate reverse bias voltage was conducted, including leakage current, analog response, fake-hit rate and charge collection efficiency.
+Reducing the sensor capacitance and expanding the depletion zone are the most intrinsic and direct consequences of applying a reverse bias on the substrate.
+After adjusting thresholds under different reverse biases and determining a set of working-points by the ITHR value scan described in Sec.
+2.2.2, a thorough sensor characterization regarding the impact of the substrate reverse bias voltage was conducted, including leakage current, analog response, fake-hit rate and charge collection efficiency.
 3.1 Leakage Current
@@
 […]
+The results, presented in Fig.
+7, indicate that the leakage current from the pixel area is small (less than 3 fA per pixel) and increases with the reverse bias applied to the substrate.
+The current observed at substrate bias 0V is a result of the potential difference between the collection node, which is held at +1V by the VReset pin, and the grounded substrate.
+This observed increase with reverse bias is consistent with the understanding that a primary component of leakage current originates from the thermal generation of electron-hole pairs within the depletion region of the pixel diode.
+As the substrate reverse bias increases, the depletion region expands, thereby enlarging the volume for carrier generation and consequently leading to a higher measured leakage current [10].
 […]
@@
 The effective pixel input capacitance of the sensor, $C_{\rm effective}$, can be determined by analyzing the analog response of the sensor under different reverse bias voltages.
-The analog output amplitude (denoted as OUT_A in Fig.
-1) is given by:
-$A_{\rm{OUT\_A}}=\frac{C_{\rm test}\cdot V_{\rm pulse}}{C_{\rm effective}}\cdot f_{\rm gain}\cdot f_{\rm sf}~{},$ (1)
-[…]
+For test purposes, the chip is equipped with several dedicated test pixels.
+The analog signal OUT_A from one of these selectable pixels can be connected to an external pin, allowing it to be monitored with an oscilloscope.
+The analog output (denoted as OUT_A in Fig.
+2) amplitude is given by:
+$A_{\rm{OUT\_A}}=\frac{Q_{\rm injected}}{C_{\rm effective}}\cdot f_{\rm gain}\cdot f_{\rm sf}~,$ (1)
+where $A_{\mathrm{OUT\_A}}$ is the amplitude of OUT_A, $Q_{\rm injected}$ is the injected charge which equals to $C_{\rm test}\cdot V_{\rm pulse}$ in the electric pulse test, $f_{\rm gain}$ is the gain factor from the front-end amplifier, and $f_{\rm sf}$ is the factor of the source follower as a buffer from the test pixel.
+$f_{\rm gain}$ varies with signal amplitude because the front-end is a non-linear amplifier.
+For the estimation of $C_{\rm effective}$, these factors were set to $f_{\rm sf}=0.7,f_{\rm gain}=20$, which are derived from post-layout SPICE simulations of the front-end circuit design.
 Two independent and complementary methods were employed in this study to measure $C_{\rm effective}$ via the relation:
@@
   • 2.
-    Method 2: In this procedure, the analog response amplitude, $A_{\rm{OUT\_A}}$, was maintained at a constant value by dynamically adjusting $V_{\rm pulse}$.
+    Method 2: In this method, the analog response amplitude, $A_{\rm{OUT\_A}}$, was maintained at a constant value by dynamically adjusting $V_{\rm pulse}$.
 The corresponding waveform measurement was repeated, and the adjusted $V_{\rm pulse}$ value (required to hold the amplitude constant) was then used in Eq.
@@
 Figure 8: Effective capacitance measured as a function of the reverse bias voltage using two methods.
-(a)Method 1: $V_{\rm pulse}$ is fixed at 160 mV, and the analog response amplitude is measured over 200 repetitions.
-(b) and (c)Method 2: The analog amplitude is fixed at 49 mV, 78 mV, respectively, by adjusting $V_{\rm pulse}$; measurements are performed at two different current threshold settings (500 pA and 650 pA, respectively).
-For both methods, the calculation of $C_{\rm effective}$ assumes $f_{\rm gain}=20$ and $f_{\rm sf}=0.7$.
+(a)Method 1: $V_{\rm pulse}$ is fixed at 160 mV.
+(b)Method 2: The analog amplitude is fixed at 49 mV by adjusting $V_{\rm pulse}$ accordingly.
+For both methods, the calculation of $C_{\rm effective}$ assumes $f_{\rm gain}=20$ and $f_{\rm sf}=0.7$ and ITHR is set at 500 pA (target working-point condition).
+The amplitude shown is the mean value from 200 repeated waveform acquisitions; the error bars represent the standard deviation (RMS) of the distribution, corresponding to the measured electronic noise.
 The two methods provide complementary insights.
 While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage.
-To examine how $f_{\rm gain}$ varies with amplitude, Method 2 was repeated at two different output levels (49 mV and 78 mV), shown in Fig.
-9 and Fig.
-9.
-The results show consistent trends with only tiny differences between the two amplitudes, confirming the reliability of the approach.
 Results for these two methods after calculation are shown in Fig.
-9.
-The input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{}{\rm V}$ to $V_{\rm sub}=-2~{}{\rm V}$, corresponding to the rapid expansion of the depletion region.
+8.
+In both methods, the input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{\rm V}$ to $V_{\rm sub}=-2~{\rm V}$, corresponding to the rapid expansion of the depletion region.
 This reduction is very important as smaller input capacitance increases the charge-voltage conversion coefficient at the input of the front-end circuit for a given charge.
-To keep the input-referred threshold constant, the VCASN was tuned to lower the OUT_A voltage baseline as the OUT_A amplitude increased with reverse bias voltage.
-This may be related to the increase of FPN and TN for reverse bias from -1 V to -5 V in Fig.
-6 6 respectively, as the same effect had been observed in a VCASN scan shown in Ref.
+As a consequence of the increased charge-voltage conversion, the amplitude of the analog signal (OUT_A) increases significantly with reverse bias.
+In Section 2.2.1, to keep the input-referred threshold constant, the VCASN was tuned at different reverse bias voltage.
+The tuning of VCASN actually lowers the effective OUT_A baseline which is relative to the voltage required to switch the transistor M9 on.
+The increase of FPN and TN in Fig.
+5 5, respectively, may be related to the tuning of VCASN because the same effect observed in a dedicated VCASN scan shown in Ref.
 [5] (Fig.
@@
 Therefore, assessing and reducing the fake-hit rate is crucial for ensuring the reliability of tracking detectors.
-To compare two distinguished readout sequence modes, the fake-hit rate for each sector $i$ in both rolling shutter (RS) and global shutter (GS) modes is calculated using a unified formula:
+JadePix-3 employs both rolling shutter (RS) and global shutter (GS) readout sequence modes.
+In the conventional rolling shutter mode, one row is activated for reading at a time [11].
+This process requires 192 ns to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$.
+While in global shutter, all 512 rows are frozen for readout after a period of acquisition time.
+To compare the two readout sequence modes, the acquisition time of GS mode is set equivalent to 98.3 $\rm{\upmu s}$ and the fake-hit rate for each sector $i$ in both modes can be calculated using a unified formula:
 $R_{i}^{\text{mode}}=\frac{N_{i}^{\text{mode}}}{N_{\text{frames}}^{\text{mode}}\cdot N_{\text{pixels}}},$ (3)
 where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector.
-In the rolling shutter mode (RS), 100,000 frames are collected ($N_{\text{frames}}^{\text{RS}}=100,000$) in the absence of external sources.
-In the global shutter mode (GS), 1,000 frames are collected ($N_{\text{frames}}^{\text{GS}}=1,000$) with an acquisition window of 98.3 $\upmu\rm s$, matching the rolling shutter readout time, to enable a comparable measurement.
-The sensitivity limit is defined as the lowest measurable fake-hit in the measurement for these two modes and is selected as $\rm 1\times 10^{-10}$ for the rolling shutter and $\rm 1\times 10^{-8}$ for the global shutter mode.
+For RS mode, the test system can readout continuously while for the GS mode the test system has to launch one readout for one frame, which limits the total number of frames that can be collected in the same amount of time.
+In the absence of external sources, a total of 100,000 frames ($N_{\text{frames}}^{\text{RS}}=100,000$) are collected in the RS mode, while 1,000 frames ($N_{\text{frames}}^{\text{GS}}=1,000$) are collected in the GS mode.
+The sensitivity limit is defined as the lowest measurable fake-hit in the measurement and is determined by the number of frames for these two modes.
+Based on Eq.
+3, this results in a sensitivity limit of $\rm 4\times 10^{-10}$ for RS mode and $\rm 4\times 10^{-8}$ for GS mode.
 As the fake-hit rate is expected to be strongly influenced by the threshold, a threshold scan under different reverse biases was performed.
 The results are shown in Fig.
-10.
-In Fig.
-10, up to 10 pixels (accounting for 0.04%) in each sector with the highest fake-hit counts are identified as hot pixels that contribute most to the false hits [20, 21].
-These hot pixels are masked.
-In comparison, values presented in Fig.
-10 are obtained without masking any pixels.
-According to these results, the difference between the two modes is insignificant above the GS sensitivity limit.
+9, which compares the fake-hit rate under two conditions: with all pixels included and with hot pixels masked.
+For the masked dataset, hot pixels were identified as those that contribute most to the false hits; specifically, up to 10 pixels (accounting for 0.04%) in each sector with the highest fake-hit counts were masked.
+According to the result, the difference between the two modes is insignificant above the GS sensitivity limit.
 Reverse bias substantially reduces the fake-hit rate compared with the unbiased state, with $V_{\rm sub}=-3~{\rm V},-4~{\rm V}$ yielding the lowest values.
-Figure 10: Fake-hit Rate measured with RS (rolling shutter) and GS (global shutter) modes versus threshold under varied reverse biases with all pixels included in Fig.
-10 and with hot pixels masked in Fig.
-10.
-In Fig.
-10, up to 10 pixels (accounting for 0.04%) for each sector, which have the largest counts of fake-hit, are considered hot pixels and masked.
-Sensitivity limits are indicated as dashed horizontal lines in every pad.
+Figure 9: Fake-hit rate as a function of threshold measured with RS (rolling shutter) and GS (global shutter) modes under various reverse biases.
+The plot compares results with all pixels included versus those with hot pixels masked.
+For the masked data, up to 10 pixels per sector (0.04%), corresponding to those with the highest fake-hit counts, were masked.
+The sensitivity limits of the measurement are indicated by the horizontal lines in each pad.
 3.4 Radioactive Source Test
 Two radioactive sources, $^{90}\rm Sr$ and $^{55}\rm Fe$, were employed to assess the sensor’s response to external ionizing radiation.
-${}^{90}\rm Sr$, a $\beta$ emitter, generates minimum ionizing particles that traverse the sensor and deposit charge over an extended region.
+$^{90}\rm Sr$, a $\beta$ emitter, generates electrons that traverse the sensor and deposit charge over an extended region.
 This distributed ionization is particularly sensitive to changes in the sensor’s depletion characteristics and charge-sharing behavior under reverse bias, thereby providing insights into the sensor’s charge collection efficiency.
-In contrast, ${}^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition.
-This characteristic makes ${}^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization.
-5.4.1 Test with ${}^{90}\rm Sr$ Source
-A ${}^{90}\rm Sr$ source was placed 2 mm above the sensor to verify JadePix-3’s response to beta rays.
-All data were collected with a 100,000-frame rolling shutter scan, ensuring equivalent collection time for every tested working-points.
-Approximately 30,000 clusters were recorded in sector 0,1,2 over the course of 50 seconds, for which the hit map is shown in Fig.
+$^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition.
+This characteristic makes $^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization, providing a signal whose magnitude is comparable to that of a Minimum Ionizing Particle (MIP) [12].
+For the source tests, an uncollimated $^{90}\rm Sr$ source was placed 2 mm above the sensor, and a $^{55}\rm Fe$ source was positioned 8 mm above it.
+For both, data were collected with a 100,000-frame rolling shutter scan.
+Due to charge sharing, the signal from a single particle interaction is often collected by several neighboring pixels, which form a cluster.
+The cluster size is defined as the number of pixels in such a group.
+Approximately 30,000 clusters were recorded in sector 0,1,2 over the course of 50 seconds, for which the hitmap from $^{90}\rm Sr$ source test is shown in Fig.
+10, where the pixels at the edge of the array were blocked partially by the PCB.
+Figure 10: Hitmap of sectors 0,1,2 from the 90Sr source test, showing the number of hits per pixel.
+A total of 27,414 clusters (8,969 in Sec.
+0, 10,053 in Sec.
+1, and 8,392 in Sec.
+2) were recorded under $\beta$-ray irradiation from the backside during a time duration of 50 s, with ITHR value fixed at 500 pA and $V_{\rm sub}$ at 0 V.
+The pixels at the edge of the array were partially obstructed by the PCB.
+A double hit is defined as an event where the same pixel registers a hit in two consecutive frames.
+In this untriggered measurement, to ensure each particle hit is counted only once, the analysis retains the first hit from such an event but excludes the subsequent hit in the consecutive frame.
+The results of cluster size analysis for both the 90Sr and 55Fe sources are presented in Fig.
+11, which plots the average cluster size as a function of threshold for different bias voltages.
+It should be noted that the average cluster size is calculated using only detected events; particles that do not produce a signal above threshold are not included in this average.
+Fig.
+11 demonstrates a clear reduction in cluster size with increasing reverse bias.
+At a threshold of 100 $e^{-}$, for instance, the average cluster size is reduced from 16-17 pixels at 0 V bias to 8-9 pixels at -6 V bias.
+This reduction is expected considering that the stronger electric field from the applied reverse bias lowers the lateral diffusion of charge carriers.
+Consequently, more charge from particle interactions is collected locally within the seed pixel, reducing the spread to its neighbors.
+As a result, fewer surrounding pixels exceed the detection threshold, leading to a smaller average cluster size.
+It is worth noting that a certain amount of charge sharing is often desirable as it improves spatial resolution [7], which is important for the CEPC application.
+The slope of cluster size as a function of the threshold also decreases significantly in both Fig.
+11 and Fig.
 11.
-Fig.
-12 displays the data analysis result.
-Due to varying solid angles relative to the fixed source, each sector inherently accumulates a different number of hits.
-However, because each working-point was recorded with the same number of frames (and thus identical detection time), and radioactive source was kept stationary throughout the measurements, the accumulated cluster count within a given sector is directly proportional to its detection efficiency.
-Therefore, for each sector, the relative detection efficiency at a given bias voltage was obtained by normalizing its accumulated cluster count by the maximum cluster count recorded for that same sector across all bias settings.
-Figure 11: Hit map of sector 0,1,2 from the ${}^{90}\rm Sr$ radioactive source test.
-Roughly 30,000 events were recorded under $\beta$-ray irradiation from the backside during a time duration of 50 s, with ITHR value fixed at 500 pA and $V_{\rm sub}$ at 0 V.
-Figure 12: ${}^{90}\rm Sr$ radioactive source response result.
-(a) Average cluster size versus the threshold.
-(b)-(d) Relative detection efficiency versus the threshold for Sector 0, 1, 2, respectively.
-The relative detection efficiency is proportional to the accumulated number of clusters detected and normalized by the maximum number of clusters measured in each sector.
-Fig.
-12 demonstrates a clear reduction in cluster size with the increase of the bias voltage, e.g., from 16-17 to 8-9 at 100 $e^{-}$ threshold from 0 V to -6 V.
-This means more charge from particle interactions is collected within the seed pixel, reducing the spread to neighboring pixels.
-Additionally, the stronger electric field lowers lateral diffusion, causing the charge to be collected more locally.
-As a result, fewer pixels surrounding the seed pixel exceed the detection threshold, leading to smaller average cluster size.
-The slope of cluster size as a function of the threshold also decreases significantly in Fig.
-12.
-This phenomenon demonstrates that the charge collection ability of the sensor is less affected by the threshold under biased conditions, which further illustrates that the charge collection is more sufficient and robust.
-Fig.
-12-12 demonstrate an enhancement in detection efficiency as reverse bias increases from 0 V to -5 V.
-This behavior can be attributed to the depletion zone expansion.
-However, the measured results also exhibit an evident efficiency drop as the bias is further increased from –5 V to –6 V.
-One possible reason for this decline is that under excessive reverse bias, the characteristics of the MOSFETs may undergo significant changes due to the body effect, leading to a reduction in signal amplification performance.
-A similar effect under high bulk bias was also observed in Ref.
-[22], where an nMOS device biased at –6 V exhibited a 260 mV threshold voltage shift and a 40% drop in drain current.
-This suggests that the analog front-end performance of JadePix-3 may be compromised under such high bias conditions, highlighting the importance of carefully selecting the operational bias range.
-5.4.2 Test with ${}^{55}\rm Fe$ Source
-A ${}^{55}\rm Fe$ source was positioned 8 mm above the sensor to verify JadePix-3’s response to X-ray.
-The data collection and processing for ${}^{55}\rm Fe$ were identical to ${}^{90}\rm Sr$, using a 100,000-frame rolling shutter scan, recording approximately 30,000 clusters over 50 seconds, ensuring a consistent collection time window.
-Figure 13: ${}^{55}\rm Fe$ radioactive source response result.
-(a) Average cluster size versus the threshold.
-(b)-(d) Relative detection efficiency versus the threshold of Sector 0, 1, 2, respectively.
-The relative detection efficiency is obtained in the same method as for the ${}^{90}\rm Sr$ source test as shown in Fig.
-12.
-As depicted in Fig.
-13, the slope of the cluster size versus threshold curve decreases significantly when $V_{\rm sub}$ increases, providing similar results to the ${}^{90}\rm Sr$ source test, which further illustrates the charge collection is more sufficient under the biased condition.
-Fig.
-13-13 shows improvement in detection efficiency as substrate reverse bias increases.
-Comparing Fig.
-13-13 internally, sector 1 exhibits a smaller efficiency drop with increasing the threshold (approximately 10%) compared to sectors 0 (20%) and 2 (25%).
-A similar behavior is also observed for the ${}^{90}\rm Sr$ source test, as shown in Fig.
-121212.
-A plausible mechanism involves the difference of the OUT_D trigger design: sector 1 uses a D-flipflop, while both sector 0 and 2 use SR-flipflops (see Tab.
-2).
-D-flipflop is edge-triggered.
-While the SR-flipflop is level-triggered or latched.
-The discrepancy may indicate that the electrical test pulses are different from the ionization signal because the thresholds are calibrated by the electrical pulse while they showed different results on source measurements.
+This is an indicator of improved charge collection efficiency.
+A steep slope (as seen at 0 V) signifies that the measured cluster size is highly sensitive to the threshold setting, which is characteristic of charge being spread by diffusion over many pixels.
+Conversely, a flatter slope indicates that the cluster size is less dependent on the threshold.
+This is a signature of efficient charge collection, where a strong electric field concentrates the majority of the charge into the seed pixel, making the resulting cluster size insensitive to threshold variations.
+Figure 11: Average cluster size as a function of threshold for different reverse bias voltages, measured with (a) the $^{90}\rm Sr$ source and (b) the $^{55}\rm Fe$ source.
 3.5 Infrared Laser Test
-Following the configuration of the laser test platform outlined in Section 3, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam.
+Following the configuration of the laser test platform outlined in Section 2.1, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam incident from the backside of the chip.
 The laser operates at a wavelength of 1064 nm.
@@
 3.5.1 Beam Position Adjustment
-To ensure optimal focusing and achieve the minimum laser spot size, which can be tuned to 3.4 $\rm{\upmu m}$, significantly smaller than the pixel size [11], at the depth of the chip’s sensitive layer, a scan was performed along the Z-direction (perpendicular to the sensor plane) to locate the beam waist.
-This scan involves measuring the number of pixel hits as a function of the laser’s Z-position, which indicates the vertical distance to the sensor surface, at two attenuation levels: 5% and 50%.
+To ensure optimal focusing and achieve the minimum laser spot size, which can be tuned to 3.4 $\rm{\upmu m}$, significantly smaller than the pixel size, at the depth of the chip’s sensitive layer, a scan was performed along the Z-direction (perpendicular to the sensor plane) to locate the beam waist, a method previously employed in the characterization of this sensor [5].
+This scan involves measuring the average number of pixel hits per frame over an acquisition of 1,000 frames as a function of the laser’s Z-position, which indicates the vertical distance to the sensor surface.
+The scan was performed at two attenuation levels of the laser: 5% and 50%.
 As shown in Fig.
-14, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm.
+12, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm.
 This position was subsequently fixed for all subsequent measurements to ensure optimal focus and minimal spot size on the chip.
-Figure 14: Number of pixel hits as a function of the laser position Z.
+Figure 12: Average number of pixel hits per frame as a function of the laser position Z.
 The upper and lower curves are for the laser power attenuation at 5% and 50%, respectively.
-Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frame numbers = 1,000.
+Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frame number = 1,000.
 To approach the center of a single pixel, a two-dimensional XY-scan along the sensor surface was performed, as illustrated in Fig.
-15.
-The cluster size during this scan was restricted to either 0 or 1 under a 75% laser tune setting, ensuring precise delineation of the pixel boundaries.
+13.
+For this scan, the laser intensity was significantly attenuated.
+This ensures that the charge deposited by a single laser pulse is small enough that at most one pixel exceeds the detection threshold, resulting in a cluster size of 0 or 1.
+This method provides a precise delineation of the pixel boundaries.
+Since the attenuated laser intensity is very close to the pixel’s threshold, this measurement is particularly sensitive to small pixel-to-pixel variations in response, which explains the observed asymmetries in Fig.
+13.
 The experimental conditions included a laser frequency of 25 kHz, a threshold of 215 $e^{-}$ for the sensor, a substrate bias voltage $V_{\rm sub}$ of 0 V, and a total of 10,000 Rolling Shutter frames.
-This scan confirmed the exact center position of a single pixel in both the X- and Y-directions.
-Figure 15: XY-scan results from a laser source.
-The plot on the left shows the hit map, the number of laser pulse events recorded at a specific (X, Y) position.
+This scan confirmed the exact center position of a single pixel in the X- and Y-directions.
+Figure 13: XY-scan results from a laser source over a selected pixel in Sector 2 (16 $\times$ 23.11 $\rm{\upmu m^{2}}$).
+The plot on the left shows the hit map, representing the number of recorded hits at each (X, Y) position.
 The upper right panel displays the hit profile along the line at Y = 9.774 mm, while the lower right panel shows the hit profile along the line at X = 20.263 mm.
@@
 3.5.2 In-Pixel Detection Efficiency
-Building on previous beam test results [13], we further probed the in-pixel detection performance by using the attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency in response to external excitation.
+Previous beam test results [6] indicated a non-uniform in-pixel detection efficiency, particularly between the long and short edges of the pixel.
+To investigate this observation, we further probed the in-pixel performance by using an attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency.
 Based on the XY-scan results as shown in Fig.
-15, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels.
+13, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels.
 These four positions, annotated in Fig.
-15, were representative of distinct in-pixel locations and were selected to investigate spatial variations in detection efficiency.
+13, were representative of distinct in-pixel locations and were selected to investigate spatial variations in detection efficiency.
 As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s detection efficiency in response to it to mimic the pathways of charged particles.
@@
 Given that 1000 laser pulse triggers are expected during this period, the detection efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses.
-Results shown in Fig.
-16 demonstrate that applying a reverse bias from 0 to –2 V had already enhanced the detection efficiency at in-pixel positions A, B, and C to nearly 100%, underscoring the performance improvement enabled by the reverse bias.
-These observations are in agreement with beam tests of other MAPS sensors [23], providing an imperative reference to further beam test results.
+The results shown in Fig.
+14 reveal distinct trends across the four in-pixel positions, providing information about the in-pixel variance.
+Position A (the pixel center) consistently exhibits the highest efficiency, while Position D (the corner) shows the lowest, reflecting diminished charge collection efficiency at the pixel boundary.
+Position B (long-edge center) and C (short-edge center) display intermediate efficiencies.
+Notably, the efficiency at Position B is higher than at Position C.
+This aligns with findings from the JadePix-3 beam test measurement [6], which also revealed a higher detection efficiency at the long-edge center compared to that at the short-edge center.
+Furthermore, the detection efficiency increases significantly with reverse bias across all positions.
+Applying a bias of just –2 V is sufficient to raise the efficiency at the center and edge positions (A, B, and C) to nearly 100%, underscoring the performance improvement enabled by the expanding depletion volume.
+It is also noteworthy that an efficiency drop is observed at Position D when the bias increases from -5 V to -6 V.
+This may be caused by a smaller lateral depletion region under such a high bias condition, highlighting the importance of carefully selecting the operational bias voltage.
 Figure 14: Detection efficiency in response to laser beam as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 13.
-Furthermore, a laser with tuning method was developed and performed in order to quantitatively characterize the sensor’s charge collection capability.
-The most straightforward way would be to measure the charge deposit in the sensor under external excitation, however, JadePix-3 employs a binary digital readout, and it lacks the analog output of the collected charge.
-Therefore, an external probe was required whose intensity can be continuously varied to infer the pixel’s charge collection sensitivity.
-By using the laser’s tunable attenuation as a proxy for the amount of injected charge, the sensor’s binary readout can be translated into a quantitative efficiency index in a method described as follows.
-The laser frequency was set to 25 Hz, and the chip was operated with 20,000 Rolling Shutter frames, equivalent to an 8-second acquisition window.
-Under these conditions, if every laser pulse generates enough charge that exceeds the detection threshold, 200 hits would be recorded by the pixel.
-The laser pulse energy was then attenuated by incrementally increasing the tune value until the number of recorded hits dropped to around 100333The actual experimental values are not exactly 100 and range between 90 and 110 due to unavoidable fluctuation..
-The tune value at this point, designated as the Detection Sensitivity Index (DSI), was recorded for each of the four marked positions.
-This DSI is defined as the laser tune value at which the number of detected hits drops to approximately 100 out of 200 possible triggers; a higher DSI indicates that the pixel remains responsive even at lower laser intensities, thus signifying more effective charge collection efficiency and thus improved detection performance.
-Figure 17: Detection Sensitivity Index (DSI) as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 15.
-The variation of the DSI with reverse bias $V_{\rm sub}$ was measured and is depicted in Fig.
-17.
-The resulting curves reveal distinct trends across the four in-pixel positions, providing more intricate information about the in-pixel variance.
-Position A, the pixel center, consistently exhibited the highest DSI, indicating peak detection efficiency, whereas Position D showed the lowest DSI, reflecting diminished sensitivity at the pixel boundary.
-Position B (long-edge center) and C (short-edge center) displayed intermediate DSI, with Position B substantially outperforming Position C.
-This aligns with findings from the JadePix-3 beam test measurement, which revealed notably higher detection efficiency at the long-edge center (Position B) compared to the short-edge center (Position C), where efficiency was significantly lower and closely resembled that of the pixel corner (Position D).
-Notably, the DSI increased with increasing reverse bias across all positions, suggesting enhanced charge collection efficiency at higher $V_{\rm sub}$ due to the expanding depletion volume.
 3.6 Optimal Reverse Bias Setting
-A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that a reverse bias of approximately –5 V offers the best overall trade-off.
+A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that, under the tuning method presented in this paper, a reverse bias of approximately –5 V offers the best overall trade-off.
 This optimal working-point not only improves detection efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
@@
 To further characterize the performance of JadePix-3, a prototype MAPS developed for the CEPC vertex detector, detailed laboratory-based measurements were performed, in particular with substrate reverse bias voltages applied.
-A comprehensive test workflow was developed and implemented for the systematic characterization of JadePix-3.
+A systematic test workflow was implemented for the characterization of JadePix-3.
 Experimental results show that the JadePix-3 sensor can work stably when a substrate reverse bias is applied.
 Moreover, the reverse bias has a considerable influence on the performance of the sensor, typically being beneficial.
-Our results reveal that applying reverse bias substantially enhances sensor performance, reducing input capacitance, enlarging the depletion region, and improving charge collection efficiency, while also lowering fake-hit rates.
-Optimal reverse bias setting is also determined within this integrated test procedure.
-This test methodology not only confirms JadePix-3’s potential for further performance improvement but also establishes a robust methodology for the evaluation and optimization of similar MAPS devices in future particle tracking applications.
-This result acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test.
-Moreover, this work will serve as the architectural and procedural foundation for the tuning, and optimization of the next iteration, the JadePix-5 sensor, thereby driving forward MAPS development to meet CEPC’s requirements.
+Our results reveal that applying reverse bias substantially enhances sensor performance, reducing input capacitance, improving charge collection efficiency, and lowering fake-hit rates, though care must be taken as performance degradation, such as a drop in efficiency, was observed at the highest tested bias of -6 V.
+An optimal reverse bias setting is also determined within this test procedure.
+This test methodology confirms JadePix-3’s potential for further performance improvement.
+These results acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test.
+Moreover, this work will serve as the architectural and procedural foundation for the tuning and optimization of the next sensor in the JadePix series, thereby driving forward MAPS development to meet CEPC’s requirements.
 Acknowledgement

### v3

@@
-Further Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector
+Further Characterization of the JadePix-3 CMOS Pixel Sensor for the CEPC Vertex Detector: in Dependence of Substrate Reverse Bias
 Abstract
@@
 JadePix-3 is a prototype Monolithic Active Pixel Sensor (MAPS) designed for the CEPC vertex detector.
-This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and detection efficiency.
+This paper presents a detailed laboratory-based characterization of the JadePix-3 sensor, focusing on the previously under-explored effects of substrate reverse bias voltage on key performance metrics: charge collection efficiency, average cluster size, and hit efficiency of laser.
 Systematic testing demonstrated that JadePix-3 operates reliably under reverse bias, exhibiting a reduced input capacitance, an expanded depletion region, enhanced charge collection efficiency, and a lower fake-hit rate.
@@
 The variation in the column pitch arises from the implementation of different digital logic blocks within the pixels.
-These blocks were designed to evaluate the performance of R-S flipflops versus D flipflops, as well as the inclusion or exclusion of a MASK configuration register, requiring adjustments to the pixel layout.
-All these differences are summarized in Tab.
+These blocks were designed to evaluate the performance of RS latches versus D flipflops, as well as the inclusion or exclusion of a MASK configuration register, aiming to minimize the pixel layout.
+The functional block diagram of the pixel is illustrated in Fig.
+1, while the key differences between the sectors are summarized in Tab.
 2.
+Figure 1: Pixel Schematic of the JadePix-3 sensor, illustrating the main components: the collection diode, the analog front-end and the digital readout logic.
+The digital logic includes configuration bits for electrical pulse test (D latch1) and pixel masking (D latch2).
+The “Hit Register” (blue dashed box) represents the different implementations (RS latch or D flipflop) across the sensor sector designs, as detailed in Tab.
+2.
+Further details on the logic operation can be found in Ref.
+[5] (Section 2.3).
+Table 2: Digital logic blocks for different sectors.
+Sector pixel layout digital logic block Remarks
+0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 RS latches
+1 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V1 D flipflop, w/o MASK
+2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 RS latches, w/o MASK
+3 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 Same as Sector 0 but with increased current for the analog front-end
 Figure 2: Schematic diagram of the cross-section of the sensor when a reverse bias is applied to the P-substrate via a P-well.
-Table 2: Analog front-ends and digital logic blocks for different sectors.
-Sector pixel layout digital logic block Remarks
-0 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 R-S flipflop
-1 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V1 D flipflop, w/o MASK
-2 16 $\times$ 23.11 $\rm{\upmu m^{2}}$ DGT_V2 R-S flipflop, w/o MASK
-3 16 $\times$ 26 $\rm{\upmu m^{2}}$ DGT_V0 Increased current for the front-end
 In the JadePix-3 sensor design, each pixel contains a P-well that allows for substrate biasing through a surface contact, as shown schematically in Fig.
-1.
+2.
 This P-well is electrically contiguous with the underlying deep P-well and the P-type epitaxial layer.
@@
 Figure 3: Circuit topology for the JadePix-3 analog front-end [5].
+This analog front-end was adapted from the ALPIDE [7] sensor, featuring low power consumption of 40 nW and minimum threshold below 100 $e^{-}$.
 The performance of MAPS is significantly influenced by the bias voltage applied to the substrate.
 This bias has a multifaceted impact: on one hand, it directly increases the size of the depletion zone and reduces the sensor’s input capacitance [8].
-On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the pixel circuitry (shown in Fig.
-2) tends to increase due to the body effect, consequently altering the circuit’s operating status.
+On the other hand, it also affects the characteristics of the integrated circuitry; specifically, as the substrate bias becomes more pronounced, the threshold of NMOS transistors within the analog front-end circuitry (M3, M5, M8 and M9 in Fig.
+3) tends to increase due to the body effect, consequently altering the circuit’s operating status.
 Therefore, adjusting the circuit’s operating point after applying substrate bias and investigating the resulting impact on these key performance metrics becomes essential.
@@
 [5], is shown in Fig.
-3.
+4.
 The JadePix-3 chip has been wire-bonded to a carrier board and then connected to an external FPGA-based DAQ system (see Fig.
-3).
+4).
 The FPGA board configures the chip parameters and transfers data to the PC via Ethernet.
@@
 To conduct laser experiment under the demand of high spatial resolution, a laser test system featuring three high-precision motor controllers is deployed (see Fig.
-3).
+4).
 The spatial position in three orthogonal directions of the laser source is regulated precisely by these three motors, each has a repeatable position resolution of 1 $\rm{\upmu m}$ and a minimum step size of 0.1 $\rm{\upmu m}$.
@@
 Central to this process is the accurate measurement of pixel thresholds, achieved using the S-curve method.
-The operational parameters of the pixels—threshold, fixed pattern noise (FPN), and temporal noise (TN)—were characterized using the standard “S-curve” method [8].
+The operational parameters of the pixels—threshold, fixed pattern noise (FPN), and temporal noise (TN)—were characterized using the standard “S-curve” method.
 This technique involves injecting a series of electrical pulses of increasing charge ($Q_{\rm injected}$) into each pixel and measuring the hit response.
@@
 The M6 MOSFET, biased by VCLIP, clamps excessively high OUT_A signals (see Fig.
-2).
+3).
 Given that the amplitude of OUT_A increases significantly with reverse bias due to a reduction in the sensor capacitance, VCLIP is set to 1.307 V.
@@
 The threshold distributions for these target working-points are presented in Fig.
-4.
+5.
 It is observed that these distributions at non-zero reverse biases are noticeably broader compared to zero bias, indicating increased pixel-to-pixel threshold variations.
 This broadening of the threshold distributions is quantified in Fig.
-5, which plots the FPN as a function of bias voltage.
-The FPN is observed to increase with reverse bias, reflecting the growing pixel-to-pixel variations.
+6, which plots the FPN as a function of bias voltage.
 The corresponding TN analysis is shown in Fig.
-5.
+6.
 An increase in TN is observed under reverse bias conditions.
@@
 The linear relationship between the mean threshold and the ITHR setting, depicted in Fig.
-6, is consistently observed across all tested reverse bias voltages.
+7, is consistently observed across all tested reverse bias voltages.
 This linearity not only confirms the consistency of the threshold tuning but also underscores that ITHR is a suitable parameter for threshold control under different reverse biases.
@@
 The results, presented in Fig.
-7, indicate that the leakage current from the pixel area is small (less than 3 fA per pixel) and increases with the reverse bias applied to the substrate.
+8, indicate that the leakage current from the pixel area is small (less than 3 fA per pixel) and increases with the reverse bias applied to the substrate.
 The current observed at substrate bias 0V is a result of the potential difference between the collection node, which is held at +1V by the VReset pin, and the grounded substrate.
@@
 The analog output (denoted as OUT_A in Fig.
-2) amplitude is given by:
+3) amplitude is given by:
 $A_{\rm{OUT\_A}}=\frac{Q_{\rm injected}}{C_{\rm effective}}\cdot f_{\rm gain}\cdot f_{\rm sf}~,$ (1)
@@
 The two methods provide complementary insights.
-While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage.
+While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor fixed via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage.
 Results for these two methods after calculation are shown in Fig.
-8.
+9.
 In both methods, the input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{\rm V}$ to $V_{\rm sub}=-2~{\rm V}$, corresponding to the rapid expansion of the depletion region.
@@
 The increase of FPN and TN in Fig.
-5 5, respectively, may be related to the tuning of VCASN because the same effect observed in a dedicated VCASN scan shown in Ref.
+6 and 6, respectively, may be related to the tuning of VCASN because the same effect observed in a dedicated VCASN scan shown in Ref.
 [5] (Fig.
@@
 JadePix-3 employs both rolling shutter (RS) and global shutter (GS) readout sequence modes.
-In the conventional rolling shutter mode, one row is activated for reading at a time [11].
+In the conventional rolling shutter mode, one row is activated for reading at a time.
 This process requires 192 ns to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$.
 While in global shutter, all 512 rows are frozen for readout after a period of acquisition time.
-To compare the two readout sequence modes, the acquisition time of GS mode is set equivalent to 98.3 $\rm{\upmu s}$ and the fake-hit rate for each sector $i$ in both modes can be calculated using a unified formula:
+To compare the two readout sequence modes, the acquisition time of the GS mode is set equivalent to 98.3 $\rm{\upmu s}$ and the fake-hit rate for each sector $i$ in both modes can be calculated using a unified formula:
 $R_{i}^{\text{mode}}=\frac{N_{i}^{\text{mode}}}{N_{\text{frames}}^{\text{mode}}\cdot N_{\text{pixels}}},$ (3)
 where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector.
-For RS mode, the test system can readout continuously while for the GS mode the test system has to launch one readout for one frame, which limits the total number of frames that can be collected in the same amount of time.
+For RS mode, the test system can readout continuously while for the GS mode the test system has to launch one readout for one frame, which limits the total number of frames that can be collected in the test.
 In the absence of external sources, a total of 100,000 frames ($N_{\text{frames}}^{\text{RS}}=100,000$) are collected in the RS mode, while 1,000 frames ($N_{\text{frames}}^{\text{GS}}=1,000$) are collected in the GS mode.
-The sensitivity limit is defined as the lowest measurable fake-hit in the measurement and is determined by the number of frames for these two modes.
+The sensitivity limit is defined as the lowest measurable fake-hit rate in the measurement and is determined by the number of frames for these two modes.
 Based on Eq.
-3, this results in a sensitivity limit of $\rm 4\times 10^{-10}$ for RS mode and $\rm 4\times 10^{-8}$ for GS mode.
+3, this results in a sensitivity limit of $\rm 4\times 10^{-10}$ for the RS mode and $\rm 4\times 10^{-8}$ for the GS mode.
 As the fake-hit rate is expected to be strongly influenced by the threshold, a threshold scan under different reverse biases was performed.
 The results are shown in Fig.
-9, which compares the fake-hit rate under two conditions: with all pixels included and with hot pixels masked.
+10, which compares the fake-hit rate under two conditions: with all pixels included and with hot pixels masked.
 For the masked dataset, hot pixels were identified as those that contribute most to the false hits; specifically, up to 10 pixels (accounting for 0.04%) in each sector with the highest fake-hit counts were masked.
@@
 $^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition.
-This characteristic makes $^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization, providing a signal whose magnitude is comparable to that of a Minimum Ionizing Particle (MIP) [12].
+This characteristic makes $^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization, providing a signal whose magnitude is comparable to that of a minimum ionizing particle (about 66 electron-hole pairs are generated per micron [11] in the 20 $\rm{\upmu m}$ epitaxial layer of the JadePix-3 prototype).
 For the source tests, an uncollimated $^{90}\rm Sr$ source was placed 2 mm above the sensor, and a $^{55}\rm Fe$ source was positioned 8 mm above it.
@@
 Approximately 30,000 clusters were recorded in sector 0,1,2 over the course of 50 seconds, for which the hitmap from $^{90}\rm Sr$ source test is shown in Fig.
-10, where the pixels at the edge of the array were blocked partially by the PCB.
+11, where the pixels at the edge of the array were blocked partially by the PCB.
 Figure 11: Hitmap of sectors 0,1,2 from the 90Sr source test, showing the number of hits per pixel.
@@
 The results of cluster size analysis for both the 90Sr and 55Fe sources are presented in Fig.
-11, which plots the average cluster size as a function of threshold for different bias voltages.
+12, which plots the average cluster size as a function of threshold for different bias voltages.
 It should be noted that the average cluster size is calculated using only detected events; particles that do not produce a signal above threshold are not included in this average.
@@
 12 and Fig.
-11.
+12.
 This is an indicator of improved charge collection efficiency.
@@
 3.5 Infrared Laser Test
-Following the configuration of the laser test platform outlined in Section 2.1, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam incident from the backside of the chip.
+Following the configuration of the laser test platform outlined in Section 2.1, a systematic investigation of the in-pixel laser hit efficiency was performed using an infrared laser beam incident from the backside of the chip.
 The laser operates at a wavelength of 1064 nm.
@@
 As shown in Fig.
-12, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm.
+13, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm.
 This position was subsequently fixed for all subsequent measurements to ensure optimal focus and minimal spot size on the chip.
@@
 To approach the center of a single pixel, a two-dimensional XY-scan along the sensor surface was performed, as illustrated in Fig.
-13.
+14.
 For this scan, the laser intensity was significantly attenuated.
@@
 Since the attenuated laser intensity is very close to the pixel’s threshold, this measurement is particularly sensitive to small pixel-to-pixel variations in response, which explains the observed asymmetries in Fig.
-13.
+14.
 The experimental conditions included a laser frequency of 25 kHz, a threshold of 215 $e^{-}$ for the sensor, a substrate bias voltage $V_{\rm sub}$ of 0 V, and a total of 10,000 Rolling Shutter frames.
@@
 Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frames = 10,000.
-3.5.2 In-Pixel Detection Efficiency
+3.5.2 In-Pixel Hit Efficiency
 Previous beam test results [6] indicated a non-uniform in-pixel detection efficiency, particularly between the long and short edges of the pixel.
-To investigate this observation, we further probed the in-pixel performance by using an attenuation-tunable infrared laser as an accessible proxy for detection efficiency and charge collection efficiency.
-Based on the XY-scan results as shown in Fig.
-13, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels.
+To investigate this observation, we further probed the in-pixel performance by using an attenuation-tunable infrared laser as an accessible proxy for hit efficiency and charge collection efficiency.
+Based on the XY-scan results shown in Fig.
+14, four distinct in-pixel positions were selected and marked: (A) the pixel center, (B) the center of the long edge, (C) the center of the short edge, and (D) the diagonal center of $2\times 2$ neighbouring pixels.
 These four positions, annotated in Fig.
-13, were representative of distinct in-pixel locations and were selected to investigate spatial variations in detection efficiency.
-As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s detection efficiency in response to it to mimic the pathways of charged particles.
-An RS scan was performed over 100,000 frames (40 seconds) under a 25 Hz laser operation for each reverse bias value across the four in-pixel positions.
-Given that 1000 laser pulse triggers are expected during this period, the detection efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses.
+14, were representative of distinct in-pixel locations and were selected to investigate spatial variations in hit efficiency.
+As the infrared laser penetrates the silicon bulk and generates electrons and holes along the way, we assessed the pixel’s hit efficiency in response to it to mimic the pathways of charged particles.
+For this laboratory measurement, we define the hit efficiency as the probability that an incident laser pulse results in a registered hit above the threshold.
+The laser intensity was kept the same as in the preceding XY-scan, but the pulse frequency was lowered to 25 Hz to ensure that a maximum of one laser pulse could arrive within a single frame.
+A 40-second Rolling Shutter scan was performed for each reverse bias value across the four in-pixel positions.
+Given that 1000 laser pulse triggers are expected during this period, the hit efficiency is calculated as the ratio of the measured hit count to 1000, thereby quantifying the pixel’s ability to detect individual laser pulses.
 The results shown in Fig.
@@
 Notably, the efficiency at Position B is higher than at Position C.
-This aligns with findings from the JadePix-3 beam test measurement [6], which also revealed a higher detection efficiency at the long-edge center compared to that at the short-edge center.
-Furthermore, the detection efficiency increases significantly with reverse bias across all positions.
+This aligns with findings from the JadePix-3 beam test measurement [6], which also revealed a higher hit efficiency at the long-edge center compared to that at the short-edge center.
+Furthermore, the hit efficiency increases significantly with reverse bias across all positions.
 Applying a bias of just –2 V is sufficient to raise the efficiency at the center and edge positions (A, B, and C) to nearly 100%, underscoring the performance improvement enabled by the expanding depletion volume.
@@
 This may be caused by a smaller lateral depletion region under such a high bias condition, highlighting the importance of carefully selecting the operational bias voltage.
-Figure 14: Detection efficiency in response to laser beam as a function of the reverse bias $V_{\rm sub}$. Measured at four in-pixel spots A, B, C, and D as marked in Fig. 13.
+Figure 15: Hit efficiency in response to laser beam as a function of the reverse bias $V_{\rm sub}$.
+Measured at four in-pixel spots A, B, C, and D as marked in Fig.
+14.
+The laser intensity was the same as that used for the XY-scan, but the frequency was reduced to 25 Hz to ensure at most one pulse per frame.
 3.6 Optimal Reverse Bias Setting
 A comprehensive survey of the JadePix-3’s performance over substrate biases from 0 to –6 V indicates that, under the tuning method presented in this paper, a reverse bias of approximately –5 V offers the best overall trade-off.
-This optimal working-point not only improves detection efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
+This optimal working-point not only improves hit efficiency and charge collection ability, but also maintains a low fake-hit rate and input capacitance.
 4 Conclusions and Outlook
@@
 These results acquired in laboratory conditions will also provide critical insights to inform subsequent beam test configurations, parameter optimization, as well as the radiation hardness test.
-Moreover, this work will serve as the architectural and procedural foundation for the tuning and optimization of the next sensor in the JadePix series, thereby driving forward MAPS development to meet CEPC’s requirements.
+Moreover, this work will serve as the architectural and procedural foundation for the tuning and optimization of the next sensor in the JadePix series, thereby driving forward MAPS development to meet CEPC’s requirements, including the ultimate 3 $\rm{\upmu m}$ spatial resolution target.
+Achieving this will require further dedicated R&D, commencing with extensive beam tests to quantify the impact of the reverse-bias settings explored in this study on the spatial resolution.
+A critical part of this future work will be to pinpoint the root causes of degradation in position resolution under different reverse bias voltages—whether stemming from the chip design, experimental setup, or analysis algorithms—in order to guide the next sensor iteration.
 Acknowledgement